OpenStudy (anonymous):

Here is one kind of similar to the last one if anyone was good at the converging/diverging problems :)

OpenStudy (anonymous):

Question #1

OpenStudy (anonymous):

this counts as algebra?

OpenStudy (anonymous):

here is the deal, if you are not familiar with these as \(n\to \infty\) anything like \(\frac{1}{n}\) or \(\frac{5}{n}\) or \(\frac{(-1)^n}{n}\) goes to zero

OpenStudy (anonymous):

since that stuff goes to zero, and you have \[\frac{(-1)^n}{n}+3\] as \(n\to \infty\) the terms get close and closer to \(3\)

OpenStudy (anonymous):

Apparently it does, and now I see thank you! You have been a great help!

OpenStudy (anonymous):

yw, glad to help this is usually done in some more advanced classes because it takes a minute to get a real definition of \[\lim_{n\to \infty}a_n=L\]